OpenStudy (megannicole016):

Consumer math question? You owe $1,240.24 on a credit card with a 13.75% APR. You decide to pay $250.00/month toward paying off the credit card. How much is the difference in the interest paid after one month if you pay at the beginning of the month compared to paying at the end of the month?

OpenStudy (megannicole016):

@pooja195 can you help?

OpenStudy (megannicole016):

@Compassionate

OpenStudy (dumbcow):

Using present value equations: Pay Beginning of month \[250 (\frac{1-v^n}{1-v}) = 1240.24\] solving for n: \[n = \frac{\ln(1-\frac{1240.24(1-v)}{250})}{\ln v}\] Pay End of Month \[250v (\frac{1-v^n}{1-v}) = 1240.24\] solving for n \[n = \frac{\ln(1-\frac{1240.24(1-v)}{250v})}{\ln v}\] where \[v = \frac{1}{1+i} = \frac{1}{1+ \frac{.1375}{12}} = 0.988671\] This gives 2 values for n Total Interest = \[\text {250n - 1240.24}\] Find difference between total interest